I have a lake with a perimeter of 3,675 ft and I would like to estimate the acerage. I understand that ft. is a mesure of length and acre is a measure of land but I thought there was a way.
thank you

Answer by checkley77(12844)

Thank you. I would call it an irregular quadrilateral with the 1st side=429ft, 2nd side=602ft, 3rd side=475 ft, and the 4th side=718ft.
Using these figures and constructing a diagonal I have 2 non-right triangles:
718 by 475 by 801
602 by 429 by 801
Using the formula:
Area=sqrt[s(s-a)(s-b)(s-c)] where s=(a+b+c)/2
s=(718+475+801)/2=1,994/2=997
area=sqrt[997(997-718)(997-475)(997-801)
area=sqrt[997*279*522*196]
area=sqrt[28,459,412,860]
area=168,699 ft^2 for one of the triangles.
s=(602+429+801)/2=1,832/2=916
area=sqrt[916(916-602)(916-429)(916-801)]
area=sqrt[916*314*487*115]
area=sqrt16,108,382,120
area=126,919 ft^2 for the other triangle.
168,699+126,919=295,618 ft^2 estimaate for the area of the lake.